OpenStudy (anonymous):

what is the answer to 6 = 2(x + 8) - 5x

OpenStudy (ivettef365):

first solve the parenthesis

OpenStudy (nurali):

6 = 2(x-8) - 5x (can also be): 2(x-8) - 5x = 6 ( distribute the 2) 2x - 16 - 5x = 6 (add the xs on a side and the absolute terms on the other) - 3x = 22 (divide by -3) x = -22/3 hope you get it

OpenStudy (anonymous):

what would it be in fraction form?

OpenStudy (nurali):

x = -22/3

OpenStudy (nurali):

if correct please click best response.